The Lopez-Portillo Company has $11.4 million in assets, 80 percent financed by debt and 20 percent financed by common stock. The interest rate on the debt is 9 percent and the par value of the stock is $10 per share. President Lopez-Portillo is considering two financing plans for an expansion to $22 million in assets.

Under Plan A, the debt-to-total-assets ratio will be maintained, but new debt will cost a whopping 12 percent! Under Plan B, only new common stock at $10 per share will be issued. The tax rate is 35 percent.

a. If EBIT is 10 percent on total assets, compute earnings per share (EPS) before the expansion and under the two alternatives. (Round your answers to 2 decimal places.)
  Earnings per Share 0 Current Plan A Plan B

b. What is the degree of financial leverage under each of the three plans? (Round your answers to 2 decimal places.)
  Degree of Financial Leverage Current Plan A Plan B

c. If stock could be sold at $20 per share due to increased expectations for the firm’s sales and earnings, what impact would this have on earnings per share for the two expansion alternatives? Compute earnings per share for each. (Round your answers to 2 decimal places.)
  Earnings per Share Plan A Plan B
Answer #1
Earning per Share-Current
A Earning Before interest and taxes(EBIT) $1.14 million (11.4*0.1)
B Amount of Debt $9.12 million (11.4*0.8)
C=B*0.09 Interest Cost $0.82 million
D=A-C Earning Before Taxes(EBT) $0.32 million
E=D*0.35 Tax $0.11 million
F=D-E Net Income $0.21 million
G Amount of Common Stock $2.28 million (11.4*0.2)
H=G/10 Number of Shares                 0.22800 million
H Number of Shares                 228,000
I=F/H Earning per Share-Current $0.91
Earning per Share-PLAN A
A Earning Before interest and taxes(EBIT) $2.20 million (22*0.1)
B Amount of Debt $17.60 million (22*0.8)
X Amount of new debt=(B-$9.12million) $8.48 million (17.6-9.12)
C Interest Cost on old debt $0.82 million
Y Interest Cost on New debt $1.02 million (8.48*0.12)
Z=C+Y Total Interest Cost $1.838 million
D=A-Z Earning Before Taxes(EBT) $0.362 million
E=D*0.35 Tax $0.127 million
F=D-E Net Income $0.235 million
G Amount of Common Stock $4.40 million (22*0.2)
H=G/10 Number of Shares                 0.44000 million
H Number of Shares                 440,000
I=F/H Earning per Share-PLAN A $0.53
Earning per Share-PLAN B
A Earning Before interest and taxes(EBIT) $2.20 million (22*0.1)
B Amount of Debt $9.12 million
C=B*0.09 Interest Cost $0.82 million
D=A-C Earning Before Taxes(EBT) $1.38 million
E=D*0.35 Tax $0.48 million
F=D-E Net Income $0.90 million
G Amount of Common Stock $12.88 million (2.28+(22-11.4)
H=G/10 Number of Shares                 1.28800 million
H Number of Shares             1,288,000
I=F/H Earning per Share-PLAN B $0.70
Earning Per Share
CURRENT $0.91
PLAN A $0.53
PLAN B $0.70
b DEGREE OF FINANCIAL LEVERAGE
EBIT EBT EBIT/EBT
FINANCIAL LEVERAGE
CURRENT $1.14 $0.32 3.571428571
PLAN A $2.20 $0.362 6.084070796
PLAN B $2.20 $1.38 1.59512761
c IMPACT ON EPS
Plan A
Amount of Common Stock $4.40 million
Number of old Shares                 0.22800 million
Amount of New Stocks $2.12 million (4.4-2.28)
Number of New Shares                      0.106 million (2.12/20)
Total Number of shares                   0.3340 million (0.228+0.106)
Earning per Share-PLAN A $0.70 (0.235/0.3340)
Plan B
Amount of Common Stock $12.88 million
Number of old Shares                 0.22800 million
Amount of New Stocks $10.60 million (12.88-2.28)
Number of New Shares                      0.530 million (10.60/20)
Total Number of shares                   0.7580 million (0.228+0.53)
Earning per Share-PLAN A $1.18 (0.9/0.7580)
Earning Per Share
PLAN A $0.70
PLAN B $1.18